OpenStudy (anonymous):

A 68-mile trip uses two gallons of gas. How many gallons would you use on a 110-mile trip? Round to the nearest tenth of a gallon.

OpenStudy (anonymous):

mpg = 68/2 = 34 gal = 110 / mpg = 110/(68/2) = 110/34 etc.